Questions

• Consider a bit streaming scenario for a video where the 

following values apply:

– The buffer size is 1 MiB

– The low watermark is set at 100KiB

– The high watermark is set at 900KiB

– The incoming data rate is 1Mbps

– The video display rate is 300 Kbps

All material presented in this course is based on the 

book by D. Dalcher and L. Brodie Successful IT 

projects


1
Expert's answer
2021-03-02T13:36:52-0500

Given,

The buffer size = 1 MiB=1000kb

The low watermark =100KiB

The high watermark =900KiB

The incoming data rate = 1Mbps=1000kbps

The video display rate =300 Kbps



Data which to be streamed =900kb- 100kb = 800 kb

Let x be the time after the streaming begins,

"1000x-300x=800"


"\Rightarrow 700x=800"


"\Rightarrow x = \frac{800}{700}"


"\Rightarrow x = \frac{8}{7}"

Hence it will take 8/7 second to fill up to the high water mark.